Use the following matrix: 0 -4-8 2 1 7 6 7 -2 A = 4 -3 -1 Find ass
Answer #1

-4 -8 2 Sol A= 4 Ghen madnx 7 6 a14 Q13 madnx. A ay SX4 asy =-& A21-4912=1, 923=-7 1 a24 G51 -3, a32 -1, as3 =7, 934- Heu a a